OpenStudy (anonymous):

If f(x) = −2x + 1, find f(−6). My answer is -7 Am I right? :)

jimthompson5910 (jim_thompson5910):

f(x) = −2x + 1 f(-6) = −2(-6) + 1 f(-6) = 12 + 1 f(-6) = 13

jimthompson5910 (jim_thompson5910):

Not sure how you got -7

jimthompson5910 (jim_thompson5910):

If you can, post your work

OpenStudy (anonymous):

I must have really messed up! Thank you for correcting me! :)

jimthompson5910 (jim_thompson5910):

its ok, just keep practicing and you'll be fine

OpenStudy (anonymous):

I found where I messed up! Thanks again! :)

jimthompson5910 (jim_thompson5910):

where did you go wrong?

OpenStudy (anonymous):

I just got confused when i wrote the problem down. I wrote -6= -2x +1 by accident... lol

jimthompson5910 (jim_thompson5910):

i gotcha

jimthompson5910 (jim_thompson5910):

that's if you wanted to find x if f(x) = -6

OpenStudy (anonymous):

Oh ok! Thanks! :)

jimthompson5910 (jim_thompson5910):

np